A garden has a perimeter of 84' and one side is 6' longer, what is the area?
Answer by josgarithmetic(39617)

List what you know:
2x+2y=84; y=6; If A is area, xy=A.

The perimeter equation uses even coefficients and constant, so is easily simplified to x+y=42, and an assignment given as y=6, substitutable into the perimeter equation as x+6=42 rendering x=36.

Now , you know dimensions x=36 and y=6, so you can computer area A.